OpenStudy (anonymous):

find the derivatives of trigonometric functions: 1. y= secx/x 2. y=xsinx+cosx 3. x^2tanx

OpenStudy (bradely):

y =secx/x Use quotient rule of derivatives dy/dx =(x(d/dx sec x)-secx(d/dx x))/x^2 =(xsecxtanx-secx)/x^2 =secx(xtanx-1)/x^2 Source: http://www.mathskey.com/question2answer/

OpenStudy (cwrw238):

2. use the product rule for x sinx

OpenStudy (cwrw238):

3 product rule again

OpenStudy (cwrw238):

2. x * d(sinx)dx + 1*sin x + d(cosx)/dx = x cos x + sin x - sin x = x cosx

OpenStudy (cwrw238):

3 x^2 d(tanx)/dx + 2x * tanx can you continue with this?